OpenStudy (anonymous):

Can someone help me? Evaluate and simplify r+s/8 for r=34 and s=14

OpenStudy (anonymous):

u mean 34 +14/8=?

OpenStudy (anonymous):

It says r+s/8, I dont know

OpenStudy (anonymous):

if v put r-34 and s=14 will result in 34 +14/8=35.75